# Collaborative Filtering Meets Spectrum Shift: Connecting User-Item Interaction with Graph-Structured Side Information

### v1

@@ Abstract.
Graph Neural Network (GNN) has demonstrated their superiority in collaborative filtering, where the user-item (U-I) interaction bipartite graph serves as the fundamental data format. However, when graph-structured side information (e.g., multimodal similarity graphs or social networks) is integrated into the U-I bipartite graph, existing graph collaborative filtering methods fall short of achieving satisfactory performance. We quantitatively analyze this problem from a spectral perspective. Recall that a bipartite graph possesses a full spectrum within the range of [-1, 1], with the highest frequency exactly achievable at $-1$ and the lowest frequency at $1$; however, we observe as more side information is incorporated, the highest frequency of the augmented adjacency matrix progressively shifts rightward. This spectrum shift phenomenon has caused previous approaches built for the full spectrum [-1, 1] to assign mismatched importance to different frequencies. To this end, we propose Spectrum Shift Correction (dubbed SSC), incorporating shifting and scaling factors to enable spectral GNNs to adapt to the shifted spectrum. Unlike previous paradigms of leveraging side information, which necessitate tailored designs for diverse data types, SSC directly connects traditional graph collaborative filtering with any graph-structured side information. Experiments on social and multimodal recommendation demonstrate the effectiveness of SSC, achieving relative improvements of up to 23% without incurring any additional computational overhead.
Spectrum shift, Collaborative filtering, Side information, Multimodal, Social
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systemsccs: Computer systems organization Neural networks
@@ 2.2. Spectral Graph Neural Networks
Since $\lambda_{\min}=-1,\lambda_{\max}=1$ holds true for any bipartite graph (see Appendix C.1 for a proof), spectral GNNs applied in collaborative filtering are typically founded on the following full-spectrum assumption:
@@
Low- and high-frequency signals are widely recognized for their effectiveness in the majority of recommendation tasks (Shen et al., 2021; Peng et al., 2022; Zhu and Zhang, 2024). Consequently, prior successful graph collaborative filtering methods often place greater emphasis on frequencies close to -1 or 1. It is readily observed that LightGCN is a special case of Eq. (2) with $\alpha_{l}=1/(L+1)$, promoting low-frequency signals as illustrated in Figure 1. In contrast, JGCF places additional emphasis on high-frequency signals. Unfortunately, the full-spectrum assumption breaks down in the context of the augmented graph introduced below. This greatly challenges the applicability of both LightGCN and JGCF. Although recent advancements also suggest a learnable polynomial (e.g., Jacobi Polynomial Bases (Wang and Zhang, 2022)), training difficulties and instability (He et al., 2020; Xu et al., 2023) arise when applying this approach to collaborative filtering. Hence, we are to endow existing spectral GNNs with a correction capability.
@@ 2.3. Graph-structured Side Information
The real-world recommendation involves perpetually accumulating data of diverse types. In collaborative filtering, graph-structured side information (Zhang et al., 2021; Jiang et al., 2024) has gained significant attention within the research community. Such information acts as a valuable supplement to interaction data, with the potential to alleviate the severe sparsity issue of bipartite graphs, particularly for cold-start users and items. We observe two types of graph-structured side information frequently discussed in social recommendation and multimodal collaborative filtering (refer to Appendix B.1 for preprocessing details), respectively.
@@ 3.2. Spectrum Shift Phenomenon
The results follow from the observation that $\tilde{A}_{+}(\kappa)$ converges to an adjacency matrix determined by $\tilde{S}$ (w.r.t. $\kappa$-rescaling) or a positive semidefinite matrix (w.r.t. $\kappa$-nearest-neighbors) as $\kappa\rightarrow+\infty$. Additional details can be found in Appendix C.2. ∎
@@ 4.1. Applications
The universal procedures for boosting spectral GNNs are detailed in Algorithm 1. Here, we introduce two specific examples.
@@ 5.1. Experiment Setup
Baselines. To validate the effectiveness of SSC, we compare it with several research lines: I) traditional collaborative filtering methods: MF-BPR (Rendle et al., 2009), LightGCN (He et al., 2020), JGCF (Guo et al., 2023); II) spectral GNNs with learnable polynomial filters: ChebyNet (Defferrard et al., 2016), JacobiConv (Wang and Zhang, 2022); III) state-of-the-art multimodal collaborative filtering methods: MMGCN (Wei et al., 2019), LATTICE (Zhang et al., 2021), BM3 (Zhou et al., 2023), FREEDOM (Zhou and Shen, 2023), MMSSL (Wei et al., 2023); IV) state-of-the-art social collaborative filtering methods: DiffNet++ (Wu et al., 2020), DSL (Wang et al., 2023), SHaRe (Jiang et al., 2024). Further details are provided in Appendix B.2.
@@ 5.3. SSC Visualization (RQ3)
Figure 3. Effect of spectrum shift correction.
In this section, we employ JGCF as the backbone to illustrate the impact of SSC on $\tilde{A}_{+}$. The shifting and scaling factors are jointly tuned under $\kappa=0,5,10,20$, with the optimal hyperparameter configuration determined based on NDCG@20 performance on the validation set. Subsequently, the spectrum of $\tilde{A}_{+}(\kappa)$ after applying spectrum shift correction is visualized in Figure 3 using green lines. While the original oracle spectrum shifts rightward, as indicated by the red lines, the adjusted spectra under various $\kappa$ values are more closely aligned within the range of $[-1,1]$. This alignment better adheres to the full-spectrum assumption, thereby allowing JGCF’s robust filtering mechanism to be effectively leveraged for collaborative filtering while incorporating side information.
@@ 6.1. Graph Collaborative Filtering
GNNs, particularly spectral GNNs, exhibit superiority in collaborative filtering due to the bipartite graph structure inherent in the interaction data. Since LightGCN (He et al., 2020) simplifies NGCF (Wang et al., 2019) by eliminating all trainable components except for the embeddings, more and more researchers have begun to recognize the significance of the spectrum (Mao et al., 2021; Peng et al., 2022). Shen et al. (2021) unified a series of classic works from a spectral perspective, highlighting the low-frequency enhancement function of LightGCN. Furthermore, the highest-frequency signals are mined in BSPM (Choi et al., 2023) and JGCF (Guo et al., 2023) through graph diffusion and Jacobi polynomial bases, respectively. However, these spectral GNNs are constrained to the bipartite interaction graph that adheres to the full-spectrum assumption. For the augmented graph examined in this paper, the underlying graph filters need to be modified to achieve optimal performance.
@@ Appendix A Oracle Spectra
Appendix B Experiment Details
B.1. Graph-structured Side Information Preprocessing
B.2. Baselines
In this paper, we consider the following baselines:
Appendix C Proofs
C.1. Full-Spectrum Assumption

### v2

@@ Abstract.
Graph Neural Networks (GNNs) have demonstrated their superiority in collaborative filtering, where the user-item (U-I) interaction bipartite graph serves as the fundamental data format. However, when graph-structured side information (e.g., multimodal similarity graphs or social networks) is integrated into the U-I bipartite graph, existing graph collaborative filtering methods fall short of achieving satisfactory performance. We quantitatively analyze this problem from a spectral perspective. Recall that a bipartite graph possesses a full spectrum within the range of [-1, 1], with the highest frequency exactly achievable at $-1$ and the lowest frequency at $1$; however, we observe as more side information is incorporated, the highest frequency of the augmented adjacency matrix progressively shifts rightward. This spectrum shift phenomenon has caused previous approaches built for the full spectrum [-1, 1] to assign mismatched importance to different frequencies. To this end, we propose Spectrum Shift Correction (dubbed SSC), incorporating shifting and scaling factors to enable spectral GNNs to adapt to the shifted spectrum. Unlike previous paradigms of leveraging side information, which necessitate tailored designs for diverse data types, SSC directly connects traditional graph collaborative filtering with any graph-structured side information. Experiments on social and multimodal recommendation demonstrate the effectiveness of SSC, achieving relative improvements of up to 23% without incurring any additional computational overhead. Our code is available at https://github.com/yhhe2004/SSC-KDD.
Spectrum Shift, Collaborative Filtering, Side Information, Multimodal, Social Network
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadadoi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systemsccs: Computer systems organization Neural networks
@@ 2.2. Spectral Graph Neural Networks
Since $\lambda_{\min}=-1,\lambda_{\max}=1$ holds true for any bipartite graph (see Appendix B.1 for a proof), spectral GNNs applied in collaborative filtering are typically founded on the following full-spectrum assumption:
@@
Low- and high-frequency signals are widely recognized for their effectiveness in the majority of recommendation tasks (Shen et al., 2021; Peng et al., 2022a; Zhu and Zhang, 2024). Consequently, prior successful graph collaborative filtering methods often place greater emphasis on frequencies close to -1 or 1. It is readily observed that LightGCN is a special case of Eq. (2) with $\alpha_{l}=1/(L+1)$, promoting low-frequency signals as illustrated in Figure 1. In contrast, JGCF places additional emphasis on high-frequency signals. Unfortunately, the full-spectrum assumption breaks down in the context of the augmented graph introduced below. This greatly challenges the applicability of both LightGCN and JGCF. Although recent advancements also suggest a learnable polynomial (e.g., Jacobi Polynomial Bases (Wang and Zhang, 2022)), training difficulties and instability (He et al., 2020; Xu et al., 2023) arise when applying this approach to collaborative filtering. Hence, we are to endow existing spectral GNNs with a correction capability.
@@ 2.3. Graph-Structured Side Information
The real-world recommendation involves perpetually accumulating data of diverse types. In collaborative filtering, graph-structured side information (Zhang et al., 2021; Jiang et al., 2024) has gained significant attention within the research community. Such information acts as a valuable supplement to interaction data, with the potential to alleviate the severe sparsity issue of bipartite graphs, particularly for cold-start users and items. We observe two types of graph-structured side information frequently discussed in social recommendation and multimodal collaborative filtering (refer to Appendix C.1 for preprocessing details), respectively.
@@ 3.2. Spectrum Shift Phenomenon
The results follow from the observation that $\tilde{A}_{+}(\kappa)$ converges to an adjacency matrix determined by $\tilde{S}$ (w.r.t. $\kappa$-rescaling) or a positive semidefinite matrix (w.r.t. $\kappa$-nearest-neighbors) as $\kappa\rightarrow+\infty$. Additional details can be found in Appendix B.2. ∎
@@ 4. Spectrum Shift Correction
max_eigen_val = 1.
min_eigen_val = 1 - power_iteration(I - A,
num_iterations=100)
mu = (min_eigen_val + max_eigen_val) / 2
delta = (max_eigen_val - min_eigen_val) / 2
Algorithm 1 Python-style algorithm for factor estimation
4.1. Simple yet Effective Method
@@
Factor estimation. The shifting and scaling factors can be estimated automatically by leveraging approximation algorithms. In the following, we illustrate how this procedure can be implemented efficiently using the classical power method. Since $\lambda_{\max}(\tilde{A})\equiv 1$, we are to estimate the minimum eigenvalue according to the Python-style pseudocode 1, wherein power_iteration is an eigenvalue algorithm that returns the largest (in absolute value) eigenvalue of the given matrix. Accordingly, we have:
[…]
and the ‘optimal’ $\mu$ and $\Delta$. Remark that an exact spectral correction may not necessarily be optimal. This is because the spectrum shift induced by the augmented graph is distributionally similar to a shifting and scaling transformation, but not completely equivalent. Indeed, the real transformation is considerably more intricate. Nonetheless, the estimated values of $\mu$ and $\Delta$ yield acceptable performance. To achieve optimal results, it is necessary to retune them in the vicinity of the estimated values.
@@ 4.2. Applications
Figure 3. Pipeline of SSC.
The universal procedures for boosting spectral GNNs are detailed in Algorithm 2 and Figure 3. Here, we introduce two specific examples.
@@ 5.1. Experiment Setup
Baselines. To validate the effectiveness of SSC, we compare it with several research lines:
I) traditional collaborative filtering methods:
II) spectral GNNs with learnable polynomial filters:
III) state-of-the-art multimodal collaborative filtering methods:
IV) state-of-the-art social collaborative filtering methods:
@@ 5.3. SSC Visualization (RQ3)
In this section, we employ JGCF as the backbone to illustrate the impact of SSC on $\tilde{A}_{+}$. The shifting and scaling factors are jointly tuned under $\kappa=0,5,10,20$, with the optimal hyperparameter configuration determined based on NDCG@20 performance on the validation set. Subsequently, the spectrum of $\tilde{A}_{+}(\kappa)$ after applying spectrum shift correction is visualized in Figure 4 using green lines. While the original oracle spectrum shifts rightward, as indicated by the red lines, the adjusted spectra under various $\kappa$ values are more closely aligned within the range of $[-1,1]$. This alignment more closely adheres to the full-spectrum assumption, enabling the backbone model’s robust filtering mechanism to be effectively leveraged for collaborative filtering while incorporating side information.
@@ 5.4. Robustness of SSC Against Noise (RQ4)
5.5. Sensitivity Analysis
Table 5. Performance under different values of $\mu$ ($\;{}^{*}$ indicates estimated value).
$\mu$ R@10 R@20 N@10 N@20
0 0.0663 0.1033 0.0359 0.0454
0.05 0.0669 0.1038 0.0361 0.0455
0.1 0.0665 0.1043 0.0360 0.0457
0.1203* 0.0663 0.1045 0.0360 0.0458
0.15 0.0652 0.1034 0.0353 0.0451
0.2 0.0643 0.1019 0.0348 0.0444
Table 6. Performance under different values of $\Delta$ ($\;{}^{*}$ indicates estimated value).
$\Delta$ R@10 R@20 N@10 N@20
1 0.0654 0.1032 0.0355 0.0452
0.9 0.0657 0.1034 0.0358 0.0454
0.8 0.0667 0.1038 0.0361 0.0456
0.8797* 0.0659 0.1042 0.0357 0.0455
0.7 0.0666 0.1031 0.0362 0.0455
0.6 0.0657 0.1006 0.0356 0.0446
We present the sensitivity experiment results on the Baby dataset in Tables 5 and 6, which also demonstrate the robustness of our rough estimation method introduced in Algorithm 1.
@@ 6.1. Graph Collaborative Filtering
GNNs, particularly spectral GNNs, exhibit superiority in collaborative filtering due to the bipartite graph structure inherent in the interaction data. Since LightGCN (He et al., 2020) simplifies NGCF (Wang et al., 2019) by eliminating all trainable components except for the embeddings, more and more researchers have begun to recognize the significance of the spectrum (Mao et al., 2021; Peng et al., 2022a). Shen et al. (2021) unified a series of classic works from a spectral perspective, highlighting the low-frequency enhancement function of LightGCN. Furthermore, the highest-frequency signals are mined in BSPM (Choi et al., 2023) and JGCF (Guo et al., 2023) through graph diffusion and Jacobi polynomial bases, respectively. However, these spectral GNNs are constrained to the bipartite interaction graph that adheres to the full-spectrum assumption. For the augmented graph examined in this paper, the underlying graph filters need to be modified to achieve optimal performance.
@@ 7. Conclusion and Future Work
8. Acknowledgments
This work was supported in part by National Natural Science Foundation of China (No. 92270119) and East China Normal University Academic Excellence Program for Future Leading Talents (Advanced Project, Grant No.2025ZY-012).
We thank all reviewers for their insightful feedback during the rebuttal period. We are especially grateful to Reviewer VXX2 for inspiring our methodology for hyperparameter estimation via the eigenvalue algorithm.
@@ Appendix A Oracle Spectra
Appendix B Proofs
B.1. Full-Spectrum Assumption
@@ B.2. Spectrum Shift
Appendix C Experiment Details
C.1. Graph-structured Side Information Preprocessing
C.2. Hyperparameter Settings
Table 7. Hyperparameters used for different backbones on each dataset.
LightGCN JGCF
$\kappa$ $\mu$ $\Delta$ $\kappa$ $\mu$ $\Delta$
Baby 10 0.05 0.75 5 0.05 0.85
Sports 5 0.1 0.7 5 0.05 0.85
Electronics 10 0.0 0.55 5 0.0 0.75
Ciao 0.75 0.15 0.4 0.75 0.0 0.55
LastFM 1.0 0.0 0.5 0.75 0.0 1.0
Table 7 summarizes the tuned hyperparameters across various dataset and model combinations.
Appendix D Additional Experiments
D.1. Other Baselines and Backbones
Table 8. Additional performance comparison for more baselines and more backbones.
Baby Ciao
R@10 R@20 N@10 N@20 R@10 R@20 N@10 N@20
GF-CF 0.0350 0.0629 0.0130 0.0202 0.0014 0.0028 0.0007 0.0012
SVD-GCN 0.0513 0.0802 0.0278 0.0352 0.0471 0.0740 0.0299 0.0378
SMORE 0.0625 0.0946 0.0344 0.0427 - - - -
STAIR 0.0674 0.1042 0.0359 0.0453 - - - -
LightGCN 0.0543 0.0851 0.0293 0.0372 0.0473 0.0770 0.0295 0.0381
   +SSC 0.0666 0.1044 0.0361 0.0458 0.0547 0.0861 0.0340 0.0431
$\blacktriangle$22.61% 22.69% 23.33% 23.20% 15.71% 11.80% 15.30% 13.06%
AFDGCF 0.0552 0.0861 0.0297 0.0377 0.0483 0.0779 0.0301 0.0387
   +SSC 0.0676 0.1037 0.0366 0.0459 0.0536 0.0851 0.0329 0.0421
$\blacktriangle$22.61% 20.48% 23.26% 21.79% 10.85% 9.15% 9.41% 8.87%
SGL 0.0550 0.0861 0.0299 0.0379 0.0475 0.0759 0.0296 0.0380
   +SSC 0.0673 0.1024 0.0363 0.0454 0.0540 0.0861 0.0336 0.0429
$\blacktriangle$22.43% 18.86% 21.66% 19.78% 13.84% 13.36% 13.52% 13.06%
SimGCL 0.0513 0.0802 0.0278 0.0352 0.0471 0.0740 0.0299 0.0378
   +SSC 0.0676 0.1048 0.0366 0.0462 0.0533 0.0852 0.0334 0.0427
$\blacktriangle$22.12% 21.21% 23.18% 22.29% 8.02% 8.52% 9.27% 9.24%
To comprehensively demonstrate the superiority of SSC over competitive baselines, we compare it with several representative methods, including direct eigenvalue decomposition (GF-CF (Shen et al., 2021), SVD-GCN (Peng et al., 2022b)) and recent state-of-the-art multimodal recommendation models (SMORE (Ong and Khong, 2025), STAIR (Xu et al., 2024)). To further showcase its generalizability, we apply SSC to several alternative backbone architectures (AFDGCF (Wu et al., 2024), SGL (Wu et al., 2021), SimGCL (Yu et al., 2022)), validating its capability to provide orthogonal improvements across different model designs. The results are shown in Table 8.
D.2. Larger datasets
Table 9. Dataset statistics.
#Users #Items #Interactions
Books 603,668 367,982 8,898,041
Yelp 186,328 94,570 2,127,869
Table 10. Performance comparison on Books.
R@10 R@20 N@10 N@20
FREEDOM OOM OOM OOM OOM
SMORE OOM OOM OOM OOM
LightGCN 0.0596 0.0878 0.0348 0.0423
    +SSC 0.0632 0.0932 0.0369 0.0449
$\blacktriangle$6.04% 6.15% 6.03% 6.15%
Table 11. Performance comparison on Yelp.
R@10 R@20 N@10 N@20
DSL 0.0674 0.1055 0.0367 0.0468
SHaRe OOM OOM OOM OOM
LightGCN 0.0645 0.1010 0.0352 0.0448
    +SSC 0.0689 0.1079 0.0379 0.0483
$\blacktriangle$6.87% 6.87% 7.69% 7.83%
We evaluate SSC on two larger datasets: Amazon Books [^1] https://github.com/enoche/MMRec (He and McAuley, 2016) for multimodal recommendation and Yelp [^2] https://business.yelp.com/data/resources/open-dataset for social recommendation. Dataset statistics are presented in Table 9, and the corresponding results are reported in Tables 10 and 11.